OpenStudy (anonymous):

find derivate of - 3/ x^3

OpenStudy (anonymous):

is this +9/ X^4

OpenStudy (anonymous):

quotient rule= (x^3)(-3)'-(x^3)'(-3) all over x^6 so this = (x^3)(0)-(3x^2)(-3) all over x^6

OpenStudy (anonymous):

the problem i get from my book is find derivate f(x)=3X^3 - (8X^3/4 DIVIDED 3) - (3 / x^3

OpenStudy (anonymous):

SO MY ANSWER IS 9X^2 - 2X^-1/4 + 9X^-4 IS THIS CORRECT

OpenStudy (anonymous):

-9x^2

OpenStudy (anonymous):

but you just pulled a part from the question, thats why i solved only for that derivative because its still -9x^2/x^6

OpenStudy (anonymous):

or 9/x^4 so yea you are right 9x^-4

OpenStudy (anonymous):

but that should still be -9x^-4

OpenStudy (anonymous):

OK SO THE 3RD TERM IS NEGATIVE THANK YOU

OpenStudy (anonymous):

is the problem -(3/x^3) or -3/x^3

OpenStudy (anonymous):

itll be positive if it becomes -(-9x^-4)=9x^-4 if there isnt a negative, it will stay -9x^-4

OpenStudy (anonymous):

-(3/x^3) THE PROBLEM

OpenStudy (anonymous):

then its negative

OpenStudy (anonymous):

because its 9x^2/x^6=9/x^4=9x^-4 then -(9x^-4)=-9x^-4

OpenStudy (anonymous):

OK THANKS I JUST THOUGHT THAT BECAUSE X^3 DE DENOMINATOR IF I BRING IT UP IT WILL BE -(3X^-3) SO THEN IT WILL BE POSITIVE

OpenStudy (anonymous):

its fine, you just have to do the quotient rule, and then at the end associate the negative with the derivative you found for the section